OpenStudy (anonymous):

Find the numbers b such that the average value of f(x) = 2 + 10x - 9x^2 on the interval [0,b] is equal to 3.

myininaya (myininaya):

do you know how to find average value of a function?

OpenStudy (anonymous):

Yeah. 1/(b-a) * integral of the function from a to b

myininaya (myininaya):

right so we have \[\frac{1}{b-0} \int\limits_{0}^{b} f(x) dx =3\]

myininaya (myininaya):

where f=2+10x-9x^2 do you know how to find an andiderivative of f?

myininaya (myininaya):

in symbols I'm asking you to compute: \[\int\limits_{}^{}(2+10-9x^2) dx\]

OpenStudy (anonymous):

2x + 5x^2 - 3x^3. Plug in b and multiply by 1/b

OpenStudy (anonymous):

Would the two answers be -1/3 or 2? I did it again and got those answers

myininaya (myininaya):

\[\frac{1}{b}(2b+5b^2-3b^3)=3 \] this is what you did?

OpenStudy (anonymous):

Yup

myininaya (myininaya):

so you have 2+5b-3b^2=3

myininaya (myininaya):

-3b^2+5b-1=0

myininaya (myininaya):

use quadratic formula

OpenStudy (anonymous):

Quadratic formula gives you b = -1/3 and 2, right?

OpenStudy (anonymous):

I must've been doing something wrong over and over again

myininaya (myininaya):

just checking if i have done my arithmetic right that is the exact same quadratic you tried to find zeros for?

myininaya (myininaya):

\[b=\frac{ -5 \pm \sqrt{5^2-4(-3)(-1)}}{2(-3)}\]

myininaya (myininaya):

is this what you got next?

OpenStudy (anonymous):

The online homework doesn't like those answers. But yea, that's what I did

myininaya (myininaya):

\[b=\frac{-5 \pm \sqrt{25-12}}{-6}\]

myininaya (myininaya):

but 13 isn't a perfect square so your answers shouldn't be so pretty

OpenStudy (anonymous):

\[5/6 +- \sqrt{13}/6\]

OpenStudy (anonymous):

I had written the quadratic formula wrong every time. Oops

myininaya (myininaya):

lol

OpenStudy (anonymous):

Got it. Thank you!

myininaya (myininaya):

http://www.wolframalpha.com/input/?i=-3b%5E2%2B5b-1%3D0 i like using this site to check my work sometimes

OpenStudy (anonymous):

Wolfram Alpha? I use that religiously lol

myininaya (myininaya):

lol

myininaya (myininaya):

so yeah the quadratic formula is \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]